OpenStudy (anonymous):

What are the x and y values in this system of equations ? 2x + 4y = 8 2x = x + 8

OpenStudy (anonymous):

correction : 2y = x + 8

OpenStudy (anonymous):

2x = x + 8 2x - x = 8 x = 8 sub x = 8 into equation one 2(8) + 4y = 8 16 + 4y = 8 4y = 8-16 4y = -16 y = -16/4 y = -4

OpenStudy (anonymous):

so , how could you solve for the other value ?

OpenStudy (anonymous):

x = 8 and y =-4 we both values.

OpenStudy (anonymous):

8 isn't one of the answer choices.

OpenStudy (anonymous):

i'm pretty sure it is, otherwise how would you solve it?